OpenStudy (anonymous):

If 2x = 5, 3y = 4, and 4z = 3, what is the value of 24xyz ? a.12 b.18 c.36 d.48 e.60

OpenStudy (anonymous):

x=5/2 y=4/3 z=3/4 24xyz =24*(5/2)*(4/3)*(3/4) 60

OpenStudy (anonymous):

2x = 5 x = 5/2 3y = 4 y = 4/3 4z = 3 z = 3/4 24(5/2)(4/3)(3/4) (120/2)(4/3)(3/4) (480/6)(3/4) 1440/24 60 I agree with madrockz :)

OpenStudy (phi):

notice that you could multiply all three equations together 2x = 5, 3y = 4, and 4z = 3 2x * 3y *4z = 5*4*3 24xyz= 60